# Towards More Robust Retrieval-Augmented Generation: Evaluating RAG Under Adversarial Poisoning Attacks

### v1

@@ Abstract
Retrieval-Augmented Generation (RAG) systems have emerged as a promising solution to mitigate LLM hallucinations and enhance their performance in knowledge-intensive domains. However, these systems are vulnerable to adversarial poisoning attacks, where malicious passages injected into retrieval databases can mislead the model into generating factually incorrect outputs. In this paper, we investigate both the retrieval and the generation components of RAG systems to understand how to enhance their robustness against such attacks. From the retrieval perspective, we analyze why and how the adversarial contexts are retrieved and assess how the quality of the retrieved passages impacts downstream generation. From a generation perspective, we evaluate whether LLMs’ advanced critical thinking and internal knowledge capabilities can be leveraged to mitigate the impact of adversarial contexts, i.e., using skeptical prompting as a self-defense mechanism. Our experiments  [^1] Code can be found at https://github.com/JinyanSu1/eval_PoisonRaG. and findings provide actionable insights into designing safer and more resilient retrieval-augmented frameworks, paving the way for their reliable deployment in real-world applications.
@@ 1 Introduction
In this paper, we explore both the retrieval and the generation aspects of RAG systems to understand how to make them more robust against adversarial poisoning attacks. Specifically, we make the following key contributions:
  • We evaluate the extent to which LLMs’ internal knowledge and critical thinking abilities can enhance their resilience against adversarial manipulation through skeptical prompting. This evaluation sheds light on the emergent self-defense capabilities of LLMs when confronted with adversarial contexts.
  • We analyze the impact of the retrieved passage’s quality on downstream generation, providing a systematic understanding of how different retrieval outcomes influence the robustness and the accuracy of RAG systems.
  • Through extensive experiments, we diagnose critical safety concerns in RAG systems and offer actionable recommendations for improving their reliability and robustness in real-world applications.
@@ 2 Related Work
Adversarial Poisoning on RAG. Pan et al. (2023) warned of the dangers of disinformation pollution with LLMs, studied several attacks and proposed simple defenses based on subsamplimng of the retrieved passages and skeptical prompting. Zhong et al. (2023); Su et al. (2024) focused only on the retrieval phase and investigated using gradient information to create adversarial contexts that are likely to be retrieved by many queries. Long et al. (2024) conducted a targeted query attack, generating adversarial passage only for the target queries while not considering the influence of retrieved adversarial passages on the LLM’s outputs, while Zou et al. (2024) focused on both the retrieval phase and the generation phase by crafting adversarial passages for specific queries and directly influencing the LLM’s generation by inducing the model to output a target answer for specific questions.
RAG with Irrelevant or Incorrect Contexts. Beyond intentionally injected adversarial contexts, RAG systems inevitably introduce irrelevant or incorrect information due to the limitations of their retrievers Yin et al. (2023). Many works Asai et al. (2023); Yan et al. (2024); Chen et al. (2024); Wang et al. (2024) showed that irrelevant or incorrect contexts could detrimentally affect model performance.
@@ 3 Peering Into Adversarial Poisoning Attacks
In RAG systems, LLMs retrieve information from external knowledge bases to generate responses. A critical vulnerability arises when the knowledge base is poisoned with misleading or malicious data, allowing adversarial contexts to influence the retrieval and ultimately compromise the model’s output. As demonstrated in PoisonedRAG Zou et al. (2024), adversarial contexts can be easily generated using a black-box approach by prompting any LLM with a targeted query-answer pair (see Figure 8 in Appendix A), without access to either the retriever or the target LLM. When tested on the target LLM, these crafted contexts achieve high retrieval success rates, exposing the susceptibility of RAG systems to poisoning attacks.
@@
At first glance, it may seem counterintuitive that adversarial passages can surpass clean, untouched passages in their retrieval rankings, especially that the adversarial passages are crafted without direct access to the retriever’s architecture or the target LLM. Figure 1 gives an example of how this happens on the NQ dataset: it shows the most relevant passage retrieved from the corpus and a crafted adversarial context, along with their similarity scores as computed by Contriever. Notably, the adversarial passage has a higher similarity score than the most relevant passage from the knowledge base, as it is explicitly tailored to the query. This pattern is further supported by an analysis of the top-10 retrieved passages for the same query (Appendix B), where 4 out of 5 passages injected are successfully retrieved and ranked highly.
@@
The previous example demonstrates how adversarial passages exploit relevance scores to achieve higher rankings, potentially because passages from the clean corpus fail to provide sufficient relevance to the query in comparison with adversarial contexts. This observation motivates us to further explore the impact of the quality of retrieved passages. Specifically, we examine factually correct passages that fall into two categories: (1) passages with lower relevance to the query than the adversarial passage, and (2) passages with a relevance level comparable to the adversarial passage. For the first category, we use the top retrieved passage from the clean knowledge base or the corpus, which we term the Untouched context. For the second category, we adopt a similar approach to crafting adversarial contexts (as shown in Figure 8), prompting an LLM to generate a passage that directly leads to the correct answer for the query. As this passage is intentionally designed to counter adversarial contexts and guide the model toward accurate outputs, we refer to it as the Guiding context.
@@ 4 Experimental Setup
Our primary evaluation measure is F1 score, which balances precision and recall to provide a comprehensive measure of performance. Additionally, we include other measures such as Precision, Recall and Abstention Rate in Appendix D.1 for reference.
@@
For the retrieval component, in addition to the adversarial context, untouched context, and guiding context discussed in the previous section, we include a Non-RAG setting as a criterion for assessing the LLM’s internal knowledge, where LLM answers queries without relying on any external contexts. For the generation component, we experiment with two types of prompts: (1) Neutral, a prompt commonly used in RAG systems that does not explicitly steer the model’s behavior, and (2) Skeptical, a prompt that explicitly encourages LLMs to critically evaluate the provided contexts and rely on their own judgment when necessary, with additional instructions such as: “Choose what you think is the correct answer if, based on your own judgment, the context is incorrect or misleading.” The exact prompts used for experiments can be found in Table 3 in Appendix C.
@@ 5.1.2 Collective Influence of Multiple Passages
In this subsection, we explore how different types of retrieved passages interact and collectively influence model performance. Specifically, we analyze the impact of strength of adversarial context pollution, the dilution effect of increasing the number of retrieved passages $k$, and the mitigating role of guiding passages. Due to space constraints, we primarily present results from the NQ dataset in the main paper, with results for other datasets available in Appendix D. These experiments highlight the complex interplay between adversarial and other passages in RAG systems.
@@ 5.2 Retrieval Analysis
In the previous section, we assumed that specific contexts had already been retrieved, allowing us to analyze the generation component with controlled contexts. In this section, we examine the retrieval component to understand the likelihood of adversarial contexts being included among the top retrieved passages. Table 1 reports the proportion of queries for which at least one of the five injected adversarial contexts is retrieved within the top-$k$ passages ($k=\{1,10\}$) across different retrievers when injecting five adversarial passages into the knowledge base. Additional results on $k=\{5,20\}$ can be found in Table 13. In HotpotQA, we observe that for all retrievers, at least 83% of the queries have at least one adversarial passage ranked higher than all untouched passages, i.e., retrieved as the top-1 passage, while in NQ, at least 47% of the queries have one adversarial passage ranked higher than all other passages in the knowledge base. These results indicate that adversarial passages frequently outrank non-adversarial contexts, significantly influencing retrieval quality. When extending the retrieval scope to the top-10 passages, we find that adversarial contexts are retrieved for more than 97% of the queries in HotpotQA and more than 83% of the queries in NQ across all retrievers. Among the evaluated retrievers, Contriever-MS appears particularly vulnerable, as adversarial passages are more frequently retrieved, thereby posing a greater risk to downstream tasks such as generation. In contrast, DPR-single demonstrates greater robustness, with adversarial contexts less likely to be retrieved prominently.
@@ 5.3 Evaluating Retrieval and Generation Together
Table 2 shows combined retrieval and generation results for the NQ dataset using top-10 retrieval. Due to space limitations, the results for HotpotQA dataset are provided in Appendix D.6, Table 14.
@@ 5.4 Further Experiments and Discussions
The purpose of skeptical prompting is not to encourage models to indiscriminately challenge all retrieved contexts. Instead, it is designed to enable models to exercise selective skepticism when a given context conflicts with their internal knowledge. When models are presented with guiding passages alone, skeptical prompting does not result in significant performance degradation compared to neutral prompting. For detailed experimental results, we refer interested readers to Appendix D.5.
@@
We include additional results on faithful prompting in Appendix D.5. Faithful prompting instructs the model to fully trust the context, which can be beneficial in specific use cases, such as domain-specific question answering or scenarios involving private knowledge bases where the external context is reliable and essential. However, for more general question answering, faithful prompting offers limited performance gains, even when the retrieved context is reliable. This might be because modern LLMs already exhibit a strong ability to follow instructions and maintain approximate fidelity to the context, explicit faithful prompting typically leads to only marginal improvements. Conversely, when the retrieved context is adversarial, imposing faithful prompting can significantly degrade performance.
@@
We found that skeptical prompting substantially reduced the LLM’s tendency to express uncertainty. Further discussion and results can be found in Appendix D.2.
@@ 7 Limitations and Future Directions
While our study provides valuable insights into the robustness of RAG systems, there are areas for further improvement and exploration. First, our experiments were conducted on commonly used question answering datasets, which may not fully represent the diverse use cases and challenges encountered in RAG systems. Expanding the analysis to more varied datasets and task domains would have an broader perspective on RAG performance under different conditions.
@@ 8 Ethical Statement
Appendix A Prompts Templates
@@
Appendix B Example of top-10 retrieved contexts
@@
Appendix C Details for Experimental Setup
@@
Precision=# Correct Answer# Correct Answer+# Incorrect $\text{Precision}=\frac{\text{\# Correct Answer}}{\text{\# Correct Answer}+\text{\# Incorrect Answer}}$
@@
Recall=# Correct Answer# Correct Answer+# Incorrect Answer+# $\text{Recall}=\frac{\text{\# Correct Answer}}{\text{\# Correct Answer}+\text{\# Incorrect Answer}+\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\# Abstain}}}$ (1)
@@
Abstention=# Abstain# Correct Answer+# Incorrect Answer+# $\text{Abstention}=\frac{\text{\# Abstain}}{\text{\# Correct Answer}+\text{\# Incorrect Answer}+\text{\# Abstain}}$ (2)
@@
Appendix D Additional experimental results
D.1 Results on other metrics
@@
D.2 Patterns of Expressing Uncertainty

### v2

@@ Abstract
Retrieval-Augmented Generation (RAG) systems have emerged as a promising solution to mitigate LLM hallucinations and enhance their performance in knowledge-intensive domains. However, these systems are vulnerable to adversarial poisoning attacks, where malicious passages injected into the retrieval corpus can mislead models into producing factually incorrect outputs. In this paper, we present a rigorously controlled empirical study of how RAG systems behave under such attacks and how their robustness can be improved. On the generation side, we introduce a structured taxonomy of context types—adversarial, untouched, and guiding—and systematically analyze their individual and combined effects on model outputs. On the retrieval side, we evaluate several retrievers to measure how easily they expose LLMs to adversarial contexts. Our findings also reveal that “skeptical prompting” can activate LLMs’ internal reasoning, enabling partial self-defense against adversarial passages, though its effectiveness depends strongly on the model’s reasoning capacity. Together, our experiments  [^1] Code is available at https://github.com/JinyanSu1/eval_PoisonRaG. and analysis provide actionable insights for designing safer and more resilient RAG systems, paving the way for more reliable real-world deployments.
@@ 1 Introduction
In this paper,we design rigorous, controlled experiments to analyze how context quality affects RAG systems, both individually and in combination-We explore both the retrieval and the generation aspects of RAG systems to understand how to make them more robust against adversarial poisoning attacks. Specifically, we make the following key contributions:
  • We introduce a structured taxonomy of context types—adversarial, untouched, and guiding—and systematically evaluate both their individual and collective influence on RAG robustness. In contrast, prior work primarily focuses on coarse-grained issues such as imperfect retrieval or knowledge conflict, which lack the granularity needed to conduct controlled experiments that isolate and analyze the effects of specific context types and their interactions.
  • We demonstrate that skeptical prompting activates LLMs’ internal reasoning capabilities, enabling them to critically assess retrieved content and defend against adversarial contexts without requiring external correction mechanisms. However, While it can serve as one practical mitigation strategy, our analysis further reveals that the effectiveness of this self-defense behavior is closely tied to the LLM’s reasoning and critical thinking capacity.
  • Through extensive experiments, we uncover several findings that raise new research questions and directions. For instance, we observe that DPR tends to embed incorrect passages farther from queries, making it less likely to retrieve adversarial contexts. We also find that skeptical prompting preserves performance on clean contexts while significantly enhancing robustness in adversarial settings.
These findings not only inform practical system design but also open up compelling avenues for future research, such as developing retriever-aware attack strategies or designing adaptive prompting mechanisms.
@@ 2 Related Work
Adversarial Poisoning on RAG. Pan et al. (2023) warned of the dangers of disinformation pollution with LLMs, studied several attacks and proposed simple defenses based on subsamplimng of the retrieved passages and skeptical prompting. Zhong et al. (2023); Su et al. (2024) focused only on the retrieval phase and investigated using gradient information to create adversarial contexts that are likely to be retrieved by many queries. Long et al. (2024) conducted a targeted query attack, generating adversarial passage only for the target queries while not considering the influence of retrieved adversarial passages on the LLM’s outputs, while Zou et al. (2024) focused on both the retrieval phase and the generation phase by crafting adversarial passages for specific queries and directly influencing the LLM’s generation by inducing the model to output a target answer for specific questions. More discussions on related work can be find in Appendix A.1.
RAG with Irrelevant or Incorrect Contexts. Beyond intentionally injected adversarial contexts, RAG systems inevitably introduce irrelevant or incorrect information due to the limitations of their retrievers Yin et al. (2023). Many works Asai et al. (2023); Yan et al. (2024); Chen et al. (2024); Wang et al. (2024) showed that irrelevant or incorrect contexts could detrimentally affect model performance. Huang et al. showed that when provided with both correct and incorrect contexts, models tend to overly rely on external information, regardless of its factual accuracy. More discussion on these related works can be find in Appendix A.1.
@@ 3 Peering Into Adversarial Poisoning Attacks
In RAG systems, LLMs retrieve information from external knowledge bases to generate responses. A critical vulnerability arises when the knowledge base is poisoned with misleading or malicious data, allowing adversarial contexts to influence the retrieval and ultimately compromise the model’s output. As demonstrated in PoisonedRAG Zou et al. (2024), adversarial contexts can be easily generated using a black-box approach by prompting any LLM with a targeted query-answer pair (see Figure 8 in Appendix B), without access to either the retriever or the target LLM. When tested on the target LLM, these crafted contexts achieve high retrieval success rates, exposing the susceptibility of RAG systems to poisoning attacks.
@@
At first glance, it may seem counterintuitive that adversarial passages can surpass clean, untouched passages in their retrieval rankings, especially that the adversarial passages are crafted without direct access to the retriever’s architecture or the target LLM. Figure 1 gives an example of how this happens on the NQ dataset: it shows the most relevant passage retrieved from the corpus and a crafted adversarial context, along with their similarity scores as computed by Contriever. Notably, the adversarial passage has a higher similarity score than the most relevant passage from the knowledge base, as it is explicitly tailored to the query. This pattern is further supported by an analysis of the top-10 retrieved passages for the same query (Appendix C), where 4 out of 5 passages injected are successfully retrieved and ranked highly.
@@
The previous example demonstrates how adversarial passages exploit relevance scores to achieve higher rankings, potentially because passages from the clean corpus fail to provide sufficient relevance to the query in comparison with adversarial contexts. This observation motivates us to further explore the impact of the quality of retrieved passages. Specifically, we examine factually correct passages that fall into two categories: (1) passages with lower relevance to the query than the adversarial passage, and (2) passages with a relevance level comparable to the adversarial passage. For the first category, we use the top retrieved passage from the clean knowledge base or the corpus, which we term the Untouched context. For the second category, we adopt a similar approach to crafting adversarial contexts (as shown in Appendix B, Figure 8), prompting an LLM to generate a passage that directly leads to the correct answer for the query. As this passage is intentionally designed to counter adversarial contexts and guide the model toward accurate outputs, we refer to it as the Guiding context.
@@ 4 Experimental Setup
Our primary evaluation measure is F1 score, which balances precision and recall to provide a comprehensive measure of performance. Additionally, we include other measures such as Precision, Recall and Abstention Rate in Appendix E.1 for reference.
@@
For the retrieval component, in addition to the adversarial context, untouched context, and guiding context discussed in the previous section, we include a Non-RAG setting as a criterion for assessing the LLM’s internal knowledge, where LLM answers queries without relying on any external contexts. For the generation component, we experiment with two types of prompts: (1) Neutral, a prompt commonly used in RAG systems that does not explicitly steer the model’s behavior, and (2) Skeptical, a prompt that explicitly encourages LLMs to critically evaluate the provided contexts and rely on their own judgment when necessary, with additional instructions such as: “Choose what you think is the correct answer if, based on your own judgment, the context is incorrect or misleading.” The exact prompts used for experiments can be found in Table 3 in Appendix D.
@@ 5.1.2 Collective Influence of Multiple Passages
In this subsection, we explore how different types of retrieved passages interact and collectively influence model performance. Specifically, we analyze the impact of strength of adversarial context pollution, the dilution effect of increasing the number of retrieved passages $k$, and the mitigating role of guiding passages. Due to space constraints, we primarily present results from the NQ dataset in the main paper, with results for other datasets available in Appendix E. These experiments highlight the complex interplay between adversarial and other passages in RAG systems.
@@ 5.2 Retrieval Analysis
[…]
@@ 5.3 Evaluating Retrieval and Generation Together
Table 2 shows combined retrieval and generation results for the NQ dataset using top-10 retrieval. Due to space limitations, the results for HotpotQA dataset are provided in Appendix E.7, Table 14.
@@ 5.4 Further Experiments and Discussions
The purpose of skeptical prompting is not to encourage models to indiscriminately challenge all retrieved contexts. Instead, it is designed to enable models to exercise selective skepticism when a given context conflicts with their internal knowledge. When models are presented with guiding passages alone, skeptical prompting does not result in significant performance degradation compared to neutral prompting. For detailed experimental results, we refer interested readers to Appendix E.6.
@@
We include additional results on faithful prompting in Appendix E.6. Faithful prompting instructs the model to fully trust the context, which can be beneficial in specific use cases, such as domain-specific question answering or scenarios involving private knowledge bases where the external context is reliable and essential. However, for more general question answering, faithful prompting offers limited performance gains, even when the retrieved context is reliable. This might be because modern LLMs already exhibit a strong ability to follow instructions and maintain approximate fidelity to the context, explicit faithful prompting typically leads to only marginal improvements. Conversely, when the retrieved context is adversarial, imposing faithful prompting can significantly degrade performance.
@@
We found that skeptical prompting substantially reduced the LLM’s tendency to express uncertainty. Further discussion and results can be found in Appendix E.2.
@@ 7 Limitations and Future Directions
While our study provides valuable insights into the robustness of RAG systems, there are areas for further improvement and exploration. First, our experiments were conducted on commonly used question answering datasets, which may not fully represent the diverse use cases and challenges encountered in RAG systems. Expanding the analysis to more varied datasets and task domains would have an broader perspective on RAG performance under different conditions. Moreover, we mainly studied adversarial poisoning attack, i.e., PoisonedRAG, benchmarking retrievers under more tailored attacks and exploring hybrid attack strategies that combine global corpus poisoning (as in PoisonedRAG) with retriever-specific optimization could be a promising avenues for future work.
@@ 8 Ethical Statement
Appendix A Additional Discussions
A.1 More Discussion of Related Works
In this section, we provide more discussions on related works.
Difference with AstuteRAG Wang et al. (2024)
While AstuteRAG explores challenges in retrieval augmentation due to imperfect or knowledge conflict, our work is distinct in the following ways: (1) We use Broader and more controlled experimental design. Specifically, We define and systematically evaluate three distinct types of contexts—Adversarial, Untouched, and Guiding—whereas AstuteRAG focuses primarily on retrieval imperfections. Our study investigates not only the effect of each type individually, but also their interactions when mixed, providing a finer-grained understanding of contextual influences in RAG. (2) We focus on emergent LLM behaviors under adversarial stress: Rather than introducing an external correction module (as in AstuteRAG), we ask whether LLMs can leverage their own knowledge and critical thinking to defend against poisoned contexts. Our skeptical prompting elicits the LLM’s internal decision-making capacity to assess the reliability of retrieved content—which is not explored in prior work. (3) We have different research goals: AstuteRAG proposes a new solution to a known issue. In contrast, we diagnose and analyze the underlying robustness dynamics of RAG using tightly controlled, interpretable experiments to inform future system design, rather than optimizing for performance on a specific dataset.
Other Adversarial Attacks
Both Xue et al. (2024) and Chaudhari et al. (2024) studied back door attack on RAG. Xue et al. (2024) propose BadRAG, which utilizes contrastive optimization to generate adversarial passages activated only by specific triggers. While Chaudhari et al. (2024) crafts a malicious poisoned document that can only be retrieved with queries containing certain trigger tokens, leading to multiple behaviors such as refusal to answer, reputation damage, privacy violations in the generation phase.
Defenses against Adversarial Attacks
Xiang et al. (2024) proposes an isolate-then-aggregate strategy: get LLM responses from each passage in isolation and then securely aggregate these isolated responses. Wei et al. identified that imperfect retrievers or noisy corpora can introduce misleading information to the retrieved contents, thus affecting the generation quality explicitly. InstructRAG is then proposed to learn the denoising process through self-synthesized rationales. Concurrent work Zhou et al. (2025) implements a cluster filtering strategy to detect potential attack patterns and then employs a self assessment process to detect malicious documents and resolve inconsistencies.
A.2 The Evolving Role of RAG
RAG was originally introduced to mitigate hallucinations in LLMs by grounding generation in external knowledge. However, as LLMs have become increasingly capable, the utility of RAG has evolved. In particular, not all queries benefit equally from retrieval, especially when the LLM already has sufficient internal knowledge to answer the question, the retrieved passages are noisy, irrelevant, or even adversarial or the domain overlaps significantly with pretraining data. Recent works Wang et al. (2024); Wei et al. ; Xue et al. (2024); Zhou et al. (2025)—including ours—suggests that blindly trusting retrieved content can be harmful, especially when retrieval quality is poor. This is especially relevant in factual QA, where accuracy is paramount, and models should ideally weigh internal knowledge against potentially untrustworthy contexts. Thus: our paper is not contradictory to the fundamental utility of RAG, but rather emphasizes the need for context-aware use of RAG, especially as LLM capabilities continue to advance. In the factual QA setting, where retrieved information may be adversarial or low-quality, encouraging LLMs to reason critically rather than blindly trust context can significantly improve robustness—without undermining the core goals of RAG.
A.3 Justification of the Dataset
We use the same dataset as used in Zou et al. (2024). Though these datasets (NQ, HotpotQA, MS-MARCO) are widely used benchmarks in the RAG literature, and serve as standardized grounds for comparative evaluation. Their use reflects realistic deployment settings and supports reproducibility. More importantly, our focus is not on evaluating LLMs’ raw knowledge memorization, but to understand how models behave when presented with retrieved contexts of varying quality, including adversarial, misleading, or irrelevant passages. Though the results in our experiments indicate that, some newer LLMs like Claude-3.5 and GPT-4o have better performance, we also observe that other state-of-the-art models such as LLaMA3-8B and LLaMA3-70B do not perform as well. This suggests that reasoning ability and robustness drive the observed performance differences, rather than because the evaluation datasets used are fairly common and likely included in the training sets of state-of-the-art models.
Appendix B Prompts Templates
@@
Appendix C Example of top-10 retrieved contexts
@@
Appendix D Details for Experimental Setup
@@
$\text{Precision}=\frac{\text{\# Correct Answer}}{\text{\# Correct Answer}+\text{\# Incorrect Answer}}$
@@
$\text{Recall}=\frac{\text{\# Correct Answer}}{\text{\# Correct Answer}+\text{\# Incorrect Answer}+\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\# Abstain}}}$ (1)
@@
$\text{Abstention}=\frac{\text{\# Abstain}}{\text{\# Correct Answer}+\text{\# Incorrect Answer}+\text{\# Abstain}}$ (2)
@@
Appendix E Additional experimental results
E.1 Results on other metrics
@@
E.2 Patterns of Expressing Uncertainty